OpenStudy (anonymous):

The equation for f(x,y) = x^4 + 6sqrt(y) -10. what is its gradient vector?

OpenStudy (nikvist):

\[grad\enspace f=\frac{\partial f}{\partial x}\vec{i}+\frac{\partial f}{\partial y}\vec{j}=4x^3\vec{i}+\frac{3}{\sqrt{y}}\vec{j}\]